# Energy Matching: Unifying Flow Matching and Energy-Based Models for Generative Modeling

### v1

@@ Abstract
Generative models often map noise to data by matching flows or scores, but these approaches become cumbersome for incorporating partial observations or additional priors. Inspired by recent advances in Wasserstein gradient flows, we propose Energy Matching, a framework that unifies flow-based approaches with the flexibility of energy-based models (EBMs). Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. Our method substantially outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ compared to $8.61$), while retaining the simulation-free training of transport-based approaches away from the data manifold. Additionally, we exploit the flexibility of our method and introduce an interaction energy for diverse mode exploration. Our approach focuses on learning a static scalar potential energy—without time conditioning, auxiliary generators, or additional networks—marking a significant departure from recent EBM methods. We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their broader adoption in generative modeling across diverse domains.
@@ 1 Introduction
Figure 1: Shown are trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching. We highlight several individual trajectories in red to illustrate their distinct behaviors. Both Action Matching and Flow Matching learn time-dependent transports and are not trained for traversing the data manifold. Conversely, EBMs and Energy Matching are driven by time-independent fields that can be iterated indefinitely, allowing trajectories to navigate across modes. While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
@@
Many generative models learn a mapping from a simple, easy-to-sample distribution (e.g., a Gaussian) to the data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) — or via iterative noising-and-denoising schemes, as in diffusion models (Ho et al., 2020; Song et al., 2021). Besides being highly effective at sample generation, diffusion and flow-based models have also been used as priors to regularize ill-posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not capture the unconditional data score explicitly, and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these smoothed manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For instance, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes the initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
@@
EBMs (Hopfield, 1982; Hinton, 2002; LeCun et al., 2006) provide an alternative approach for approximating the data distribution by learning a scalar-valued function $E(x)$ that specifies an unnormalized density $p(x)\propto\exp\left(-E(x)\right)$. Rather than explicitly mapping noise samples onto the data manifold, EBMs assign low energies to regions of high data concentration and high energy elsewhere. This defines a Boltzmann distribution, from which one can sample by, for instance, Langevin sampling. In doing so, EBMs explicitly retain the likelihood information in $E(x)$. This likelihood information can then be leveraged in conditional generation (e.g., to solve inverse problems), possibly together with additional priors by simply adding their energy terms (Du and Mordatch, 2019). Moreover, direct examination of local curvature on the data manifold—allows the computation of local intrinsic dimension (LID) (an important proxy for data complexity)—whereas diffusion models can only approximate such curvature in the proximity of noise samples.
@@
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks (Guo et al., 2023; Zhang et al., 2024), thereby requiring significantly higher parameter counts and training complexity. Although multi-network approaches can improve sample quality, they inflate parameter counts by redundantly learning the same features across multiple networks, compromising the shared statistical strength of a unified feature set and the flexibility of time-independent models.
{mdframed}[hidealllines=true,backgroundcolor=blue!5]
@@
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1. When samples lie far from the data manifold, they are transported efficiently toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on CIFAR-10 generation (FID $3.97$ vs. $8.61$) and compares favorably to flow-matching and diffusion models, without auxiliary generators or time-dependent EBM ensembles.
Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models.
@@ 2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium.
@@ 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
Depending on the approach used to compute the OT coupling, one may find an OT map or not. Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map. In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs
$\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)\quad\text{with}\quad\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}.$
In our experiments, we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023), and we used the POT solver (Flamary et al., 2021).
Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation
$x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)},$
which is a point along the geodesic (Ambrosio et al., 2008). The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$. For this, we define the loss:
@@ 2.1.3 Dual objective
To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3). We define a sampling time $\operatorname{\tau_{s}}$. In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2. The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$. We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$. The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2. Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives. Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyper-parameters for each dataset are defined in Appendix C.
@@
1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from pretrained model (Algorithm 1)
@@ 3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix C.
@@ 3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion + one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate on CIFAR-10, reporting FID in Table 1. Our method outperforms state-of-the-art EBMs reducing FID by $50\%$.
Table 1: FID score comparison for unconditional CIFAR-10 generation (lower is better). Unless otherwise specified, we use results for solvers that most closely match our setup (300 fixed-step Euler–Heun (Butcher, 2016)). indicates reproduced methods, while all other entries reflect the best reported results.
@@
Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54
EGC (Guo et al., 2023) 5.36 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45
Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68
@@
ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07
CLEL-large (38M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35
Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
@@ 3.2 Inverse problems
In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w$ is Gaussian noise. In this setting, the posterior distribution of $x$ given $y$ is
$p(x|y)\propto\underbrace{\exp\left(-\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. By taking the negative logarithm of (7), the maximum a posteriori estimate of $x$ given $y$ can be obtained by minimizing
$U_{\theta}(x;y)=\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}+E_{\theta}(x).$ (8)
Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update.
Suppose now we want to recover two images from a masked image, while additionally encouraging diverse reconstructions. EBMs readily allow us to do so by introducing an interaction energy
$W(x_{1},x_{2})=\frac{\norm{B(x_{1}-x_{2})}^{2}}{\sigma^{2}},$
where $B$ is a matrix with ones in the region of interest and zeros elsewhere, thus focusing the diversity within that particular fragment of the image, and $\sigma$ is a hyper-parameter controlling the strength of the interaction. Specifically, we define
$p(x_{1},x_{2}|y)\propto p(x_{1}|y)\cdot p(x_{2}|y)\cdot\exp\left(W(x_{1},x_{2})\right),$
giving high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$. This encourages exploring the edges of the posterior rather than just its modes. With suitable $W$, samples shift towards rare events without needing many draws. To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task. From a partially observed (masked) face, we aim to reconstruct two distinct high-fidelity completions:
@@
On the left is the masked face. On the right are the reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones, zeros elsewhere), and the measurement matrix $A$ appears in the dotted blue square (zeros inside, ones outside). The interaction yields a wider range of completions yet preserves fidelity. We detail the inverse problems generation algorithm in Algorithm 3.
1:for $m=1$ to $M$ do
3:end for
5:for $n=0$ to $N-1$ do
7:         $U_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}\leftarrow\frac{\varepsilon^{(n)}}{\zeta^{2}}\bigl{\lVert}y-A\bigl{(}x_{m}^{(n)}\bigr{)}\bigr{\rVert}^{2}+V_{\theta}\bigl{(}x_{m}^{(n)}\bigr{)}+\frac{\varepsilon^{(n)}}{\sigma^{2}}\sum_{k\neq m}W\bigl{(}x_{m}^{(n)},x_{k}^{(n)}\bigr{)}$
10:    end for
11:end for
12:return $\{x_{m}^{(N)}\}_{m=1}^{M}$triangleright Final set of samples

### v3

@@ Abstract
The most widely used generative models map noise and data distributions by matching flows or scores. However, they struggle to incorporate partial observations and additional priors—something energy-based models (EBMs) handle elegantly by simply adding corresponding scalar energy terms. We address this issue by proposing Energy Matching, a framework that endows flow-based approaches with the flexibility of EBMs. Far from the data manifold, samples move along curl-free, optimal transport paths from noise to data. As they approach the data manifold, an entropic energy term guides the system into a Boltzmann equilibrium distribution, explicitly capturing the underlying likelihood structure of the data. We parameterize this dynamic with a single time-independent scalar field, which serves as both a powerful generator and a flexible prior for effective regularization of inverse problems. Our method substantially outperforms existing EBMs on CIFAR-10 and ImageNet generation in terms of fidelity, while retaining simulation-free training of transport-based approaches away from the data manifold. Furthermore, we leverage the method’s flexibility to introduce an interaction energy that supports diverse mode exploration, which we demonstrate in a controlled protein-generation setting. Our approach focuses on learning a scalar potential energy—without time-conditioning, auxiliary generators, or additional networks—which marks a significant departure from recent EBM methods. We believe that this simplified framework significantly advances EBMs capabilities and paves the way for their wider adoption in generative modeling across diverse domains.
@@ 1 Introduction
Figure 1: Trajectories (green lines) of samples traveling from a noise distribution (black dots; here, a Gaussian mixture model) to a data distribution (blue dots; here, two moons as in (Tong et al., 2023)) under four different methods: Action Matching (Neklyudov et al., 2023), Flow Matching (OT-CFM) (Tong et al., 2023), EBMs trained via contrastive divergence (Hinton, 2002), and our proposed Energy Matching. We highlight several individual trajectories in red to illustrate their distinct behaviors. Both Action Matching and Flow Matching learn time-dependent transports and are not trained for traversing the data manifold. Conversely, EBMs and Energy Matching are driven by time-independent fields that can be iterated indefinitely, allowing trajectories to navigate across modes. While samples from EBMs often require additional steps to equilibrate (see, e.g., the visible mode collapses that slow down sampling from the data manifold), Energy Matching directs samples toward the data distribution in “straight” paths, without hindering the exploration of the data manifold.
@@
Many generative models learn to map from a simple, easy-to-sample distribution, such as a Gaussian, to a desired data distribution. They do so by approximating the optimal transport (OT) map—such as in flow matching (Lipman et al., 2023; Liu et al., 2023; Tong et al., 2023) - or through iterative noising and denoising schemes, such as in diffusion models (Ho et al., 2020; Song et al., 2021). In addition to being highly effective in sample generation, diffusion- and flow-based models have also been used as priors to regularize poorly posed inverse problems (Chung et al., 2023; Mardani et al., 2024; Ben-Hamu et al., 2024). However, these models do not explicitly capture the unconditional data score and instead model the score of smoothed manifolds at different noise levels. The measurement likelihood, on the other hand, is not tractable on these noised manifolds. As a result, existing approaches repeatedly shuttle between noised and data distributions, leading to crude approximations of complex, intractable terms Daras et al. (2024). For example, DPS (Chung et al., 2023) approximates an intractable integral using a single sample. More recently, D-Flow (Ben-Hamu et al., 2024) optimizes initial noise by differentiating through the simulated trajectory. To the best of our knowledge, these models lack a direct way to navigate the data manifold in search of the optimal solution without repeatedly transitioning between noised and data distributions.
@@
EBMs (Hopfield, 1982; Hinton, 2002; LeCun et al., 2006) provide an alternative approach for approximating the data distribution by learning a scalar-valued function $E(x)$ that specifies an unnormalized density $p(x)\propto\exp\left(-E(x)\right)$. Rather than explicitly mapping noise samples onto the data manifold, EBMs assign low energies to regions of high data concentration and high energy elsewhere. This defines a Boltzmann distribution from which one can sample, for example, via Langevin sampling. In doing so, EBMs explicitly retain the likelihood information in $E(x)$. This likelihood information can then be used in conditional generation (e.g., to solve inverse problems), possibly together with additional priors simply by adding their energy terms (Du and Mordatch, 2019). Moreover, direct examination of local curvature on the data manifold—allows the computation of local intrinsic dimension (LID) (an important proxy for data complexity)—whereas diffusion models can only approximate such curvature in the proximity of noise samples.
@@
Despite the theoretical elegance of using a single, time-independent scalar energy, practical EBMs have historically suffered from poor generation quality, falling short of the performance of diffusion or flow matching models. Traditional methods (Song and Kingma, 2021) for training EBMs, such as contrastive divergence via Markov chain Monte Carlo (MCMC) or local score-based approaches (Song and Ermon, 2019), often fail to adequately explore the energy landscape in high-dimensional spaces, leading to instabilities and mode collapse. Consequently, many methods resort to time-conditioned ensembles (Gao et al., 2021), hierarchical latent ensembles (Cui and Han, 2024), or combine EBMs with separate generator networks trained in cooperation (Guo et al., 2023; Zhang et al., 2024; Yoon et al., 2024), thereby requiring significantly higher parameter counts and training complexity. {mdframed}[hidealllines=true,backgroundcolor=blue!5]
@@
In this work, we propose Energy Matching, a two-regime training strategy that combines the strengths of EBMs and flow matching; see Figure 1.
When samples lie far from the data manifold, they are efficiently transported toward the data. Once near the data manifold, the flow transitions into Langevin steps governed by an internal energy component, enabling precise exploration of the Boltzmann-like density well around the data distribution. This straightforward approach produces a time-independent scalar energy field whose gradient both accelerates sampling and shapes the final density well—via a contrastive objective that directly learns the score at the data manifold—yet remains efficient and stable to train. Empirically, our method significantly outperforms existing EBMs on both CIFAR-10 and ImageNet generation in terms of fidelity, and compares favorably to flow-matching and diffusion models—without auxiliary generators or time-dependent EBM ensembles.
Our proposed process complements the advantages of flow matching with an explicit likelihood modeling, enabling traversal of the data manifold without repeatedly shuffling between noise and data distributions. This simplifies both inverse problem solving and controlled generations under a prior. In addition, to encourage diverse exploration of the data distribution, we showcase how repulsive interaction energies can be easily and effectively incorporated, with an application to conditional protein generation. Finally, we also showcase how analyzing the learned energy reveals insight on the LID of the data with fewer approximations than diffusion models. [^1] The code will be publicly released upon acceptance.
@@ 2.1 Training objectives
In practice, we balance the two objectives by initially training $V_{\theta}$ exclusively with the optimal-transport-like objective ($\varepsilon=0$, see Section 2.1.1), ensuring a stable and consistent generation of high-quality negative samples for the contrastive phase. Subsequently, we jointly optimize both the transport-based and contrastive divergence objectives, progressively increasing the effective temperature to $\varepsilon=\varepsilon_{\max}$ as samples approach equilibrium (see Section 2.1.2).
@@ 2.1.1 Flow-like objective $\operatorname{\mathcal{L}_{\mathrm{OT}}}$
Depending on the approach used to compute the OT coupling, one may find an OT map or not. Similarly, if the set of noise samples has cardinality smaller or larger than $B$, the OT coupling will not be expressed by a map. In this case, one can adapt our algorithm by defining a threshold $\pi_{\mathrm{th}}$ and consider all the pairs $\left(x_{0}^{(j)},x_{\mathrm{data}}^{(i)}\right)$ with $\gamma^{\ast}(x_{0}^{(j)},x_{\mathrm{data}}^{(i)})>\pi_{\mathrm{th}}$. In our experiments, we used the POT solver (Flamary et al., 2021) and we did not notice benefits in a sample size other than $B$, in line with the observations in (Tong et al., 2023).
Then, for each data point $x_{\mathrm{data}}^{(i)}$ we define the interpolation $x_{t}^{(i)}=(1-t)T(x_{\mathrm{data}}^{(i)})+tx_{\mathrm{data}}^{(i)}$, which is a point along the geodesic (Ambrosio et al., 2008). The velocity of each $x_{t}^{(i)}$ is $x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$ (i.e., the samples move from the noise to the data distribution at constant speed) and, in this regime, we would like to have $-\nabla_{x}V_{\theta}(x_{t}^{(i)})\approx x_{\mathrm{data}}^{(i)}-T(x_{\mathrm{data}}^{(i)})$. For this, we define the loss:
@@ 2.1.3 Dual objective
To balance the flow-like loss $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ with the contrastive divergence loss $\operatorname{\mathcal{L}_{\mathrm{CD}}}$, we adopt the linear temperature schedule (3). We define a sampling time $\operatorname{\tau_{s}}$. In fact, although convergence to the equilibrium distribution is guaranteed only as $\operatorname{\tau_{s}}\to\infty$, we empirically observe that sample quality (measured with Fréchet inception distance (FID)) plateaus by $\operatorname{\tau_{s}}=3.0$ on CIFAR-10; see Section A.2. The time discretization determines the temperature samples $\varepsilon^{(m)}\vcentcolon=\varepsilon(m\Delta t)$. We define a dataset-specific hyperparameter $\lambda_{\mathrm{CD}}$ to scale $\operatorname{\mathcal{L}_{\mathrm{CD}}}$. The resulting algorithm is described in detail in Algorithm 1 and Algorithm 2. Since Algorithm 2 benefits from high-quality negatives, we begin with Algorithm 1 (and, thus, with $\operatorname{\mathcal{L}_{\mathrm{OT}}}$ only) to ensure sufficient mixing of noise-initialized negatives. Section A.1 discusses how the landscape of $V_{\theta}$ evolves across these two phases. Hyper-parameters for each dataset are defined in Appendix D.
@@
1:$\theta\leftarrow\theta_{\text{pretrained}}$triangleright Initialize from the warmed-up model (Algorithm 1)
@@ 3 Applications
In this section, we demonstrate the effectiveness and versatility of our proposed Energy Matching approach across three applications: (i) unconditional generation (ii) inverse problems, and (iii) LID estimation. The model architecture and all the training details are reported in Appendix D.
@@ 3.1 Unconditional generation
We compare four classes of generative models: (1) Diffusion models, which deliver state-of-the-art quality but typically require many sampling steps; (2) Flow-based methods, which learn OT paths for more efficient sampling with fewer steps; (3) EBMs, which directly model the log-density as a scalar field, offering flexibility for inverse problems and constraints but sometimes at the expense of sample quality; and (4) Ensembles (Diffusion with one or many EBMs), which combine diffusion’s robust sampling with elements of EBM flexibility but can become large and complex to train. Our approach, Energy Matching, offers a simple (a single time-independent scalar field) yet powerful EBM-based framework. We evaluate our approach on CIFAR-10 (Krizhevsky et al., 2009) and ImageNet32x32 (Deng et al., 2009; Chrabaszcz et al., 2017) datasets, reporting FID scores in Table 1 and Table 2. Our method surpasses state-of-the-art EBMs, achieving a $50\%$ reduction in FID.
[…]
. Learning Unnormalized Data Likelihood Learning Transport/Score Along Noised Trajectories Ensembles: Diffusion + (one or many) EBMs Diffusion Models Hierarchical EBM Diffusion (Cui and Han, 2024) 8.93 DDPM (Ho et al., 2020) 6.54 EGC (Guo et al., 2023) 5.36 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 3.45 Cooperative DRL (40M params) (Zhu et al., 2024) 4.31 NCSN++ (107M params, 1000 steps) (Song et al., 2021) 2.45 Cooperative DRL-large (145M params) (Zhu et al., 2024) 3.68 Energy-based Models Flow-based Models ImprovedCD (Du et al., 2021) 25.1 Action Matching (Neklyudov et al., 2023) 10.07 CLEL-large (32M params) (Lee et al., 2023) 8.61 Flow-matching (Lipman et al., 2023) 6.35 Energy Matching (50M params, Ours) 3.97 OT-CFM (37M params) (Tong et al., 2023) 4.11
[…]
@@
Cooperative DRL (40M params) (Zhu et al., 2024) 9.35 DDPM++ (62M params, 1000 steps) (Kim et al., 2021) 8.42
@@
ImprovedCD (Du et al., 2021) 32.48 Flow-matching (Lipman et al., 2023) (196M params) 5.02
CLEL-base (Lee et al., 2023) (7M params) 22.16
CLEL-large (Lee et al., 2023) (32M params) 15.47
Energy Matching (50M params, Ours) 7.67
@@ 3.2 Inverse problems
In many practical applications, we are interested in recovering some data $x$ from noisy measurements $y$ generated by an operator $A$, $y=A(x)+w$, where $w\sim\mathcal{N}\!\bigl{(}0,\sqrt{2}\,\zeta I\bigr{)}$. In this setting, the posterior distribution of $x$ given $y$ is
$p(x|y)\propto\underbrace{\exp\left(-\frac{1}{\zeta^{2}}\|y-A(x)\|^{2}\right)}_{\displaystyle\propto p(y|x)}\underbrace{\exp\left(-E_{\theta}(x)\right)}_{\displaystyle\propto p(x)},$ (7)
where $E_{\theta}(x)$ is an energy function which one can learn from the data, an EBM. Because we want to sample $x$ given a measurement $y$, this reconstruction task is often referred to as an inverse problem. Here, $\|y-Ax\|^{2}$ encodes the measurement fidelity with $\zeta$ controlling the balance between this fidelity term and the prior. We obtain the prior term $E_{\theta}(x)=\frac{V_{\theta}(x)}{\varepsilon_{\max}}$ by training $V_{\theta}(x)$ via Energy Matching. Samples from this posterior can be drawn by starting from a random sample $x^{(0)}\sim\mathcal{N}(0,I)$ and following a Langevin update. We detail the algorithm for generating solutions to inverse problems in Algorithm 3 of Appendix B. We demonstrate our model’s capabilities qualitatively through a controlled inpainting task and quantitatively via a protein inverse design benchmark.
Controlled inpainting.
Suppose we want to recover two images from a masked image while encouraging diverse reconstructions. EBMs allow this by introducing an additional interaction energy, $W(x_{1},x_{2})=-\frac{\|B(x_{1}-x_{2})\|^{2}}{\sigma^{2}}$, where $B$ has ones in the region of interest (focusing diversity there) and zeros elsewhere, and $\sigma$ is a hyper-parameter controlling the interaction’s strength. Specifically, we define $p(x_{1},x_{2}\mid y)\propto p(x_{1}\mid y)\,p(x_{2}\mid y)\,\exp\bigl(-W(x_{1},x_{2})\bigr{missing})$, which gives high probability to pairs $(x_{1},x_{2})$ that lie far apart in the specified region $B$. This encourages exploring the edges of the posterior rather than just its modes, and with suitable $W$, samples shift toward rare events without needing many draws. To illustrate the interaction term’s advantages for diverse reconstruction, we apply our method to a CelebA (Liu et al., 2015) $64\times 64$ inpainting task. As shown in Figure 2, we start from a partially observed (masked) face and aim to reconstruct two distinct high-fidelity completions.
@@
Figure 2: Controlled inpainting for diverse reconstructions. On the left is the masked face. On the right are two reconstructions: the top pair without the interaction term and the bottom pair with it. The interaction term applies in the solid red square (where $B$ has ones), and the measurement matrix $A$ is the dotted blue square (zeros inside, ones outside). By encouraging $x_{1}$ and $x_{2}$ to differ in the target region, the interaction yields a wider range of completions while preserving fidelity.
Protein inverse design.
In Figure 3, we demonstrate our method’s performance on the inverse design problem of generating Adeno-Associated Virus (AAV) capsid protein segments (Bryant et al., 2021). Given a desired functional property (fitness)—here defined as the predicted viral packaging efficiency normalized between 0 and 1—the goal is to design novel protein sequences satisfying this target condition. Beyond achieving high fitness, practical inverse design requires generating diverse candidate sequences to ensure robustness in the downstream experimental validation (Jain et al., 2022). We evaluate on two benchmark splits (medium and hard), which correspond to subsets of the original AAV dataset differing in baseline fitness distributions and required mutational distance from known high-performing variants (Kirjner et al., 2024). Leveraging the latent-space representation of VLGPO (Bogensperger et al., 2025), we employ our Energy-Matching Langevin sampler with an inference-time tunable repulsion term, allowing explicit control over the diversity of the designed proteins. This enables a flexible trade-off between fitness and diversity, resulting in high fitness scores alongside substantially improved sequence diversity. See Section B.2 for experimental details and dataset descriptions.
[…]
Figure 3: Fitness–diversity trade-off for protein inverse design on the AAV Medium (left) and Hard (right) benchmarks. We compare our Energy Matching method (blue), with diversity explicitly controlled by a repulsion strength parameter ($\propto\frac{1}{\sigma^{2}}$), against leading flow-based (purple), score-based (orange), and other non-likelihood methods (black). Fitness measures how well generated sequences satisfy the target property (predicted viral packaging efficiency), while diversity quantifies the average Levenshtein distance between sequences in each generated batch.